# DGNS: Deformable Gaussian Splatting and Dynamic Neural Surface for Monocular Dynamic 3D Reconstruction

### v2

@@ Abstract
Dynamic scene reconstruction from monocular video is critical for real-world applications. This paper tackles the dual challenges of dynamic novel-view synthesis and 3D geometry reconstruction by introducing a hybrid framework —Deformable Gaussian Splatting and Dynamic Neural Surfaces (DGNS), in which both modules can leverage each other for both tasks. During training, depth maps generated by the deformable Gaussian splatting module guide the ray sampling for faster processing and provide depth supervision within the dynamic neural surface module to improve geometry reconstruction. Simultaneously, the dynamic neural surface directs the distribution of Gaussian primitives around the surface, enhancing rendering quality. To further refine depth supervision, we introduce a depth-filtering process on depth maps derived from Gaussian rasterization. Extensive experiments on public datasets demonstrate that DGNS achieves state-of-the-art performance in both novel-view synthesis and 3D reconstruction. https://github.com/Benzlxs/DGNS
1 Introduction
@@
Most scenes in our world are dynamic in nature and achieving dynamic scene reconstruction from monocular video can empower robots with strong perception capabilities, essential to many real-world applications. Dynamic scene reconstruction primarily involves two key tasks: dynamic novel-view generation and 3D reconstruction. In the field of dynamic novel-view synthesis, several approaches [26, 27, 43, 15] have extended neural radiance fields (NeRF)[37] by incorporating feature grid planes or implicit deformation fields. Another line of work[60, 24, 61] models dynamic scenes using explicit Gaussian representations, such as 3D Gaussian Splatting (3DGS) [23]. While these approaches have achieved promising visual quality, they struggle to accurately recover the 3D geometry of dynamic scenes. For 3D geometry reconstruction in dynamic scenes, some methods [3] have combined deformation fields with implicit surface representations—specifically, SDF in canonical space. Other approaches [29, 4] utilize deformable 3DGS to reconstruct dynamic surfaces by introducing strong regularization to ensure that Gaussian primitives adhere to the surface. However, these methods encounter difficulties in producing high-fidelity novel-view synthesis. As illustrated in Fig. 1, current methods tend to excel either in dynamic novel-view synthesis (top-right area) or in dynamic 3D reconstruction (bottom-left area), but balancing these two tasks remains an unresolved challenge in dynamic scene reconstruction.
@@
Figure 1: Comparison between different methods, the higher PSNR, the better. The smaller CD, the better. Methods that represent the best of both should be located in the top-left area.
Motivated by the observation that robust geometry guidance can enhance, rather than diminish, rendering quality in static scene modeling [52, 64], we propose a hybrid representation that combines deformable Gaussian and dynamic neural surfaces for dynamic scene reconstruction from monocular video. In this framework, the deformable Gaussian splatting (DGS) module is optimized primarily for appearance reconstruction, while the dynamic neural surfaces (DNS) module focuses on geometry reconstruction. The depth maps generated by the DGS module guide the ray sampling process and provide supervision within the DNS module, while the signed distance function (SDF) learned in the DNS module informs the distribution of Gaussian primitives around the surface. To mitigate the noise in depth supervision for the DNS module, we employ Gaussian rasterization to render two types of depth maps—$\alpha$-blended and median depth maps—and introduce a filtering process to create an accurate depth map for supervising depth in the DNS module. As reconstruction from monocular video is a highly under-constrained optimization problem, we introduce the normal supervision for both modules. We conducted extensive experiments on two public datasets, where our method outperformed existing approaches on both novel-view synthesis and 3D reconstruction tasks. Our primary contributions are as follows:
@@
  • We propose a hybrid representation of deformable Gaussian and dynamic neural surfaces, marking a pioneering advancement in monocular dynamic 3D reconstruction. Our approach surpasses state-of-the-art performance in both novel-view synthesis and geometry reconstruction.
@@
  • We introduce a simple yet effective depth-filtering process for depth maps generated through Gaussian rasterization, enhancing the accuracy of depth supervision.
  • We perform extensive comparative experiments on public datasets for both novel-view synthesis and geometry reconstruction, demonstrating the advantages of our method over existing approaches.
2 Related work
@@
View synthesis for dynamic scene: View synthesis for dynamic scenes is both challenging and crucial for 3D modeling. NeRF [37] has shown impressive capabilities in generating high-fidelity novel views for static scenes by using Multi-Layer Perceptrons (MLPs) to model the radiance field, which is then rendered into pixel colors through neural volumetric techniques. Extensions of NeRF to dynamic scenes[26, 27, 43, 15, 47, 31] use time-conditioned latent codes and explicit deformation fields to capture temporal variations. However, the reliance on NeRFs’ extensive point sampling along each ray and the computational demands of MLPs limit their scalability for dynamic scenes. To address these limitations, studies have introduced techniques such as hash encoding [38, 48], explicit voxel grids [54, 12, 10, 14], and feature grid planes [7, 5, 13, 45], which have accelerated training and improved performance in handling dynamic scenes. Another line of research explores geometric primitive rasterization using point clouds [1, 62, 55], offering computational efficiency and flexibility, though these methods often encounter challenges with discontinuities and outliers. More recently, 3DGS [23] has emerged as a promising approach, leveraging anisotropic 3D Gaussians as rendering primitives. These Gaussians are depth-ordered and projected onto a 2D plane through alpha-blending, enabling high-quality real-time rendering. Various approaches have extended 3DGS to dynamic scenes [60, 24, 61, 21, 33]. For example,  [33] introduced dynamic 3DGS by iteratively optimizing the Gaussians per frame, while  [53, 61, 21] used deformation fields to model temporal changes in Gaussian distributions. Despite these advancements, such techniques remain primarily focused on novel-view synthesis and often struggle to capture scene geometry accurately, resulting in limitations in high-quality surface extraction.
@@
Dynamic surface reconstruction: Reconstructing dynamic surfaces from monocular video is essential for applications such as intelligent robotics and virtual reality. Traditional approaches often depend on predefined object templates [68, 6, 20] or temporal tracking [16, 11, 67]. With advances in neural implicit 3D representations [37, 40], methods like LASR [56] and ViSER [57] reconstruct articulated shapes using differentiable rendering techniques [30], while BANMo [58] and PPR [59] apply NeRF to dynamic scenes, and SDFFlow [35] models dynamic motion by estimating derivatives of the SDF value. Other approaches utilize RGB-D data to incorporate depth information, improving supervision for dynamic object modeling. Examples include SobolevFusion [46], OcclusionFusion [28], NDR [3], and DynamicFusion [39]. Recently, 3DGS has been integrated into several methods to enhance optimization speed and robustness. Examples include MoSca [25], Dg-mesh [29], MoGS [34], DynaSurfGS [4], and Shape-of-Motion [50]. Specifically, Dg-mesh [29] introduces Gaussian-mesh anchoring to ensure Gaussians are evenly distributed, tracking mesh vertices over time and producing high-quality meshes using a differential Poisson solver [42]. Shape-of-Motion [50] employs data-driven priors, such as monocular depth maps and 2D tracks, to constrain Gaussian motion, while DynaSurfGS [4] combines Gaussian features from 4D neural voxels with planar-based splatting for high-quality rendering and surface reconstruction. Despite progress, a fidelity gap persists due to the explicit regularization of Gaussian primitives, which restricts rendering quality relative to D-3DGS. Our approach addresses this with surface-aware density control of Gaussian, improving 3D reconstruction while maintaining high fidelity.
3 Preliminary
@@
Figure 2: The framework of DGNS consists of two primary modules: the top module, DNS, for 3D reconstruction, and the bottom module, DGS, for view synthesis. There are three key interactions between these modules. The orange arrows represent the information flow from DGS to DNS, including efficient ray-sampling and filtered depth supervision. The blue arrow illustrates the information flow from DNS to DGS, implementing surface-aware density control.
3.1 Deformable 3DGS
@@
3DGS [23] uses Gaussian primitives to represent a static scene, achieving high-quality rendering fidelity. Each Gaussian primitive is defined with a center ($x_{i}$), a covariance matrix $\Sigma_{i}$, an opacity $\sigma_{i}$, and spherical harmonics coefficients $h_{i}$, i.e. $G_{i}=\{{x_{i}},\Sigma_{i},\sigma_{i},h_{i}\}$. When rendering novel-view images, the 3D Gaussian primitives are projected onto the 2D image plane and then combined using $\alpha$-blending through a tile-based differentiable rasterizer. The color C(p) of a pixel p is computed as follows:
${C(p)=\sum_{i\in N}c_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j}^{2D})},\alpha_{i}=\sigma_{i}e^{-\frac{1}{2}(p-\boldsymbol{x}_{i})^{T}\Sigma{\prime}_{i}(p-\boldsymbol{x}_{i})}$
where $\Sigma{\prime}_{i}$ is the 2D projection of the 3D Gaussian’s covariance matrix. The rasterization process accumulates each Gaussian contribution efficiently, enabling high-quality rendering. However, standard 3DGS is limited to static scenes and cannot model temporal changes.
@@
To extend 3DGS to dynamic scenes, Deformable 3DGS (D-3DGS) [61] incorporates a deformation field that models time-dependent changes in position, rotation, and scale. Each Gaussian in canonical space is dynamically transformed by applying offsets calculated through a deformation MLP network $F_{\theta}$, i.e. $(\delta\mathbf{x},\delta\mathbf{r},\delta\mathbf{s})=F_{\theta}(\gamma(\mathbf{x}),\gamma(t)),$ $F_{\theta}(.)$, where $\gamma(\cdot)$ denotes positional encoding, $\mathbf{x}$ is the Gaussian’s canonical position, and $t$ is the current time step. The deformed Gaussian at time $t$ is expressed as: $G(\mathbf{x}+\delta\mathbf{x},\mathbf{r}+\delta\mathbf{r},\mathbf{s}+\delta\mathbf{s},\sigma).$
The rasterization pipeline remains differentiable, allowing gradients to backpropagate through the Gaussian parameters and the deformation network during optimization. This framework handles both temporal consistency and fine-grained motion, achieving high-quality rendering.
3.2 Dynamic Neural SDF
@@
where $d(x)$ is the SDF value at a point $x$. Opacity used in 3DGS or NeRF can be derived from SDF value using a logistic function [49]. Neural Dynamic Reconstruction (NDR) [3] extends SDF to model dynamic objects by incorporating a deformation field. The deformation field provides a homeomorphic (continuous and bijective) mapping (.$\mathcal{H}(.,t)$: $\mathbb{R}^{3}\rightarrow\mathbb{R}^{3}$, which maps $x_{o}$ of deformable observation space at time $t$ to its corresponding point $x_{c}$ in a canonical 3D space, where the SDF is defined independently of time or motion. This formulation ensures that any point on the dynamic surface is described accurately by setting $d(x_{c})=0$. The Dynamic Neural Surface can be defined by
$ds=\{x\in\mathbb{R}^{3}\mid\mathcal{F}(x_{c})=\mathcal{F}(\mathcal{H}(x_{o},t))=0\}\vspace{-0.6em}$ (3)
@@
The deformation field $\mathcal{H}(.)$ is designed strictly invertible, allowing a point $x_{o}$ in canonical space to map back to any observed frame $t$ via the inverse transformation $H^{-1}$. Invertibility of $\mathcal{H}$ enforces a cycle-consistent constraint across frames, which is an important regularization for modeling dynamic scenes [51].
4 Method
@@
Current approaches for dynamic scene reconstruction typically excel in either novel-view synthesis or 3D geometry reconstruction, but not both. Our framework, illustrated in Fig. 2, introduces a hybrid representation that combines Deformable Gaussian Splatting and Dynamic Neural Surfaces modules, with both components jointly optimized to enhance performance across both tasks. The details of each module are presented in the following sections.
4.1 DNS for dynamic surface reconstruction
@@
The DNS module utilizes the deformation field defined in Eq.3 to map the dynamic observation space to canonical space. Unlike other dynamic SDF methods[3], our approach incorporates ray-sampling with depth proposals and depth supervision from the DGS module. Additionally, our method introduces normal supervision using surface normals from a foundation model.
@@
Efficient ray-sampling: Ray-sampling can be computationally intensive without prior information [37], as exhaustive sampling is needed to approximate a pixel’s color from an unknown density distribution accurately. Efficiency can be improved by selectively sampling in non-empty regions of the scene. NeRF’s hierarchical volume sampling scheme uses a coarse model to approximate a density distribution, guiding the sampling of the fine model and increasing computational efficiency. In addition to hierarchical sampling, two other major sampling schemes are proposal-based [2] and occupancy-based [63, 19]. Proposal-based methods, like Mip-NeRF 360 [2], replace the coarse model with a compact proposal model that produces only density rather than both density and color. Occupancy-based methods, such as PlenOctree [63], effectively filter out points with a low density, avoiding unnecessary sampling. While these coarse-to-fine strategies [37, 49, 44, 2] improve efficiency, they often require costly rendering processes. In our framework, we use the depth map generated by the DGS branch to eliminate unnecessary queries in empty or occluded regions, thereby speeding up the ray-sampling process. The DGS depth map provides proximity to the surface, constraining the sampling range. Specifically, the $\alpha$-blending depth map $d^{\alpha}$ defines these sampling boundaries, and $d^{\alpha}$ is calculated as follows:
@@
where $N$ represents the count of 3D Gaussians encountered, $d_{i}$ is the distance from the $i$-th Gaussian to the camera, and $\alpha_{i}$ denotes opacity. The sampling process starts by emitting a ray from the camera center, $\vec{o}$, along a direction $\vec{v}$. The ray-sampling points in observation space are taken near $\vec{o}+d^{\alpha}\cdot\vec{v}$, with the range adjusted based on the SDF values $s$ calculated at ray-sampling points, $d=\mathcal{F}(\mathcal{H}(\vec{o}+d^{\alpha}\cdot\vec{v},t))$ , where $\mathcal{F}$ is to predict the SDF value for a point in observation space. The sampling interval along the ray ($\vec{o}$, $\vec{v}$) in observation space is from $\vec{o}+(d^{\alpha}-s|d|)\cdot\vec{v}$ to $\vec{o}+(d^{\alpha}+s|d|)\cdot\vec{v}$, in which the $s$ is a scaling factor that adjusts based on the predicted SDF value.
Figure 3: Visualization of the depth map in 3D space. The rightmost is the RGB image, images from left to right are 3D point clouds projected from $\alpha$-blending depth, median depth, and filtered dpeth.
@@
Filtered points are on the surface with SDF loss, as follows:
@@
Apart from depth cues, we add the normal supervision to relieve the under-constrained optimization problem. The monocular normal foundation models, such as Marigold [22, 36], are used to generate the pseudo normal map $\bar{{N}}$ Similar to MonoSDF [47], we use the volume rendering method to calculate the normal of ray $\mathbf{r}$, denoted as $\hat{{N}}$ , which is a weighted sum of each ray-sampling point normal, $\nabla_{\mathbf{x_{o}}}\mathcal{F}(\mathcal{H}(x_{o},t)))$ , on the ray $\mathbf{r}$. The consistency between the volume rendered normal map $\hat{{N}}$ and the predicted monocular normal map $\bar{{N}}$ is imposed with angular and $L1$ loss [47, 65], as follows

### v3

@@ Abstract.
Dynamic scene reconstruction from monocular video is essential for real-world applications. We introduce DGNS, a hybrid framework integrating Deformable Gaussian Splatting and Dynamic Neural Surfaces, effectively addressing dynamic novel-view synthesis and 3D geometry reconstruction simultaneously. During training, depth maps generated by the deformable Gaussian splatting module guide the ray sampling for faster processing and provide depth supervision within the dynamic neural surface module to improve geometry reconstruction. Conversely, the dynamic neural surface directs the distribution of Gaussian primitives around the surface, enhancing rendering quality. In addition, we propose a depth-filtering approach to further refine depth supervision. Extensive experiments conducted on public datasets demonstrate that DGNS achieves state-of-the-art performance in 3D reconstruction, along with competitive results in novel-view synthesis [^1] https://benzlxs.github.io/dgns_project; Jie Hong is the corresponding author..
3D Gaussian Splatting, dynamic scene reconstruction, 3D reconstruction
conference: ACM multimedia; Oct 27–31, 2025; Dublin, Irelandjournalyear: 2025copyright: ccconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia ($M$ ’25), October 27–31, 2025, Dublin, Irelanddoi: 10.1145/3746027.3755446isbn: 979-8-4007-2035-2/2025/10ccs: Computing methodologies Computer vision representations
1. Introduction
@@
Most scenes in our world are dynamic, and achieving dynamic scene reconstruction from monocular video can empower robots or intelligent agents with strong perception capabilities, essential for many real-world applications. Dynamic scene reconstruction primarily involves two key tasks: dynamic novel-view generation and 3D reconstruction. In the field of dynamic novel-view synthesis, several approaches (Li et al., 2022, 2021; Pumarola et al., 2021; Gao et al., 2021) have extended neural radiance fields (NeRF) (Mildenhall et al., 2021) by incorporating feature grid planes or implicit deformation fields. Another line of work (Yang et al., 2023a; Kratimenos et al., 2023; Yang et al., 2024) models dynamic scenes using explicit Gaussian representations, such as 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023). While these approaches have achieved promising visual quality, they struggle to accurately recover the 3D geometry of dynamic scenes. For 3D geometry reconstruction in dynamic scenes, some methods (Cai et al., 2022) have combined deformation fields with implicit surface representations—specifically, signed distance function (SDF) in canonical space. Other methods (Liu et al., 2024; Cai et al., 2024) utilize deformable 3DGS to reconstruct dynamic surfaces by introducing strong regularization to ensure Gaussian primitives adhere to the surface. However, these methods encounter difficulties in producing high-fidelity novel-view synthesis. As illustrated in Fig. 1, current methods tend to excel either in dynamic novel-view synthesis (top-right area) or in 3D reconstruction (bottom-left area), but balancing these two tasks remains an unresolved challenge in dynamic scene reconstruction.
@@
Figure 1. Performance comparison between different methods on the Dg-mesh dataset. The higher PSNR, the better. The smaller CD, the better. Methods representing the best of both should be in the top-left area.
[…]
@@
  • We propose a novel hybrid representation combining deformable Gaussian splatting and dynamic neural surfaces, achieving state-of-the-art geometry reconstruction and competitive novel-view synthesis results.
@@
  • We introduce an effective depth-filtering method to enhance depth supervision from Gaussian rasterization.
  • Extensive experiments on public datasets validate the superior performance of our approach over existing methods.
2. Related work
2.1. View Synthesis for Dynamic Scene
@@
View synthesis for dynamic scenes is both challenging and crucial for 3D modeling. NeRF (Mildenhall et al., 2021) has shown impressive capabilities in generating high-fidelity novel views for static scenes by using Multi-Layer Perceptrons (MLPs) to model the radiance field, which is then rendered into pixel colors through neural volumetric techniques. Extensions of NeRF to dynamic scenes(Li et al., 2021; Pumarola et al., 2021; Gao et al., 2021; Li et al., 2022; Xiao et al., 2025; Tian et al., 2023; Liu et al., 2023) use time-conditioned latent codes and explicit deformation fields to capture temporal variations. However, the reliance on NeRFs’ extensive point sampling along each ray and the computational demands of MLPs limit their scalability for dynamic scenes. To address these limitations, studies have introduced techniques such as hash encoding (Müller et al., 2022; Wang et al., 2024a), explicit voxel grids (Fridovich-Keil et al., 2022; Fang et al., 2022; Xu et al., 2023; Gan et al., 2023; Guo et al., 2024a), and feature grid planes (Chen et al., 2022; Cao and Johnson, 2023; Fridovich-Keil et al., 2023; Shao et al., 2023), which have accelerated training and improved performance in handling dynamic scenes. Another line of research explores geometric primitive rasterization using point clouds (Yifan et al., 2019; Aliev et al., 2020; Xu et al., 2022), offering computational efficiency and flexibility, though these methods often encounter challenges with discontinuities and outliers. More recently, 3DGS (Kerbl et al., 2023) has emerged as a promising approach, leveraging anisotropic 3D Gaussians as rendering primitives. These Gaussians are depth-sorted and alpha-blended onto a 2D plane, enabling high-quality real-time rendering. Various approaches have extended 3DGS to dynamic scenes (Yang et al., 2023a; Kratimenos et al., 2023; Yang et al., 2024; Katsumata et al., 2023; Tong et al., 2025; Luiten et al., 2023; Guo et al., 2024b). For example,  (Luiten et al., 2023) introduced dynamic 3DGS by iteratively optimizing the Gaussians per frame, while D3DGS (Yang et al., 2024) used deformation fields to model temporal changes in Gaussian distributions. Despite these advancements, such techniques remain primarily focused on novel-view synthesis and often struggle to capture scene geometry accurately, resulting in limitations in high-quality surface extraction.
2.2. Dynamic Surface Reconstruction
@@
Reconstructing dynamic surfaces from monocular video is essential for applications such as intelligent robotics and virtual reality. Traditional approaches often depend on predefined object templates (Zuffi et al., 2017; Casillas-Perez et al., 2021; Kairanda et al., 2022) or temporal tracking (Zollhöfer et al., 2018; Grassal et al., 2022; Feng et al., 2021). With advances in neural implicit 3D representations (Park et al., 2019; Mildenhall et al., 2021), methods like LASR (Yang et al., 2021a) and ViSER (Yang et al., 2021b) reconstruct articulated shapes using differentiable rendering techniques (Liu et al., 2019), while BANMo (Yang et al., 2022) and PPR (Yang et al., 2023b) apply NeRF to dynamic scenes, and SDFFlow (Mao et al., 2024) models dynamic motion by estimating derivatives of the SDF value. Other approaches utilize RGB-D data to incorporate depth information, improving supervision for dynamic object modeling. Examples include SobolevFusion (Slavcheva et al., 2018), OcclusionFusion (Lin et al., 2022), NDR (Cai et al., 2022), and DynamicFusion (Newcombe et al., 2015). Recently, 3DGS has been integrated into several methods to enhance optimization speed and robustness. Examples include MoSca (Lei et al., 2024), Dg-mesh (Liu et al., 2024), MoGS (Ma et al., 2024), DynaSurfGS (Cai et al., 2024), and Shape-of-Motion (Wang et al., 2024b). Specifically, Dg-mesh (Liu et al., 2024) introduces Gaussian-mesh anchoring to ensure Gaussians are evenly distributed, tracking mesh vertices over time, and producing high-quality meshes using a differential Poisson solver (Peng et al., 2021). Shape-of-Motion (Wang et al., 2024b) employs data-driven priors, such as monocular depth maps and 2D tracks, to constrain Gaussian motion, while DynaSurfGS (Cai et al., 2024) combines Gaussian features from 4D neural voxels with planar-based splatting for high-quality rendering and surface reconstruction. Despite progress, a fidelity gap persists due to the explicit regularization of Gaussian primitives, which restricts rendering quality relative to D3DGS (Yang et al., 2024). Our approach addresses this with surface-aware density control of Gaussian, improving 3D reconstruction while maintaining high fidelity.
3. Preliminary
@@
Figure 2. The framework of DGNS consists of two primary modules: the top module, DNS, for 3D reconstruction, and the bottom module, DGS, for view synthesis. There are three key interactions between these modules. The orange arrows represent the information flow from DGS to DNS, including efficient ray-sampling and filtered depth supervision. The green arrow illustrates the information flow from DNS to DGS, implementing surface-aware density control.
3.1. Deformable 3DGS
@@
3DGS (Kerbl et al., 2023) uses Gaussian primitives to represent a static scene, achieving high-quality rendering fidelity. Each Gaussian primitive is defined with a center ($x_{i}$), a covariance matrix $\Sigma_{i}$, an opacity $\sigma_{i}$, and spherical harmonics coefficients $h_{i}$, i.e. $G_{i}=\{{x_{i}},\Sigma_{i},\sigma_{i},h_{i}\}$. When rendering novel-view images, 3D Gaussian primitives are projected onto the 2D image plane and combined using $\alpha$-blending through a tile-based differentiable rasterizer. The color C(p) of a pixel p is computed as follows:
(1) $C(p)$ $=\sum_{i\in N}c_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j}^{2D}),$
$\alpha_{i}$ $=\sigma_{i}e^{-\frac{1}{2}(p-\boldsymbol{x}_{i})^{T}\Sigma{\prime}_{i}(p-\boldsymbol{x}_{i})}$
@@
where $\Sigma{\prime}_{i}$ is the 2D projection of the 3D Gaussian’s covariance matrix. The rasterization process accumulates each Gaussian contribution efficiently, enabling high-quality rendering. However, standard 3DGS is limited to static scenes and cannot model temporal changes. To extend 3DGS to dynamic scenes, D3DGS (Yang et al., 2024) incorporates a deformation field that models time-dependent changes in position, rotation, and scale. Each Gaussian in canonical space is dynamically transformed by applying offsets calculated through a deformation MLP network $F_{\theta}$, i.e., $(\delta\mathbf{x},\delta\mathbf{r},\delta\mathbf{s})=F_{\theta}(\gamma(\mathbf{x}),\gamma(t)),$ $F_{\theta}(.)$, where $\gamma(\cdot)$ denotes positional encoding, $\mathbf{x}$ is the Gaussian’s canonical position, and $t$ is the current time step. The deformed Gaussian at time $t$ is expressed as: $G(\mathbf{x}+\delta\mathbf{x},\mathbf{r}+\delta\mathbf{r},\mathbf{s}+\delta\mathbf{s},\sigma).$ The rasterization pipeline remains differentiable, allowing gradients to backpropagate through the Gaussian parameters and the deformation network during optimization. This framework handles both temporal consistency and fine-grained motion, achieving high-quality rendering.
3.2. Dynamic Neural SDF
@@
where $d(x)$ is the SDF value at a point $x$. Opacity used in 3DGS or NeRF can be derived from SDF value using a logistic function (Wang et al., 2021). Neural Dynamic Reconstruction (NDR) (Cai et al., 2022) extends SDF to model dynamic objects by incorporating a deformation field. The deformation field provides a homeomorphic (continuous and bijective) mapping $\mathcal{H}(.,t)$: $\mathbb{R}^{3}\rightarrow\mathbb{R}^{3}$, which maps $x_{o}$ of deformable observation space at time $t$ to its corresponding point $x_{c}$ in a canonical 3D space, where the SDF is defined independently of time or motion. This formulation ensures that any point on the dynamic surface is described accurately by setting $d(x_{c})=0$. The Dynamic Neural Surface can be defined by:
(3) $ds=\{x\in\mathbb{R}^{3}\mid\mathcal{F}(x_{c})=\mathcal{F}(\mathcal{H}(x_{o},t))=0\}$
@@
where the deformation field $\mathcal{H}(.)$ is designed strictly invertible, allowing a point $x_{o}$ in canonical space to map back to any observed frame $t$ via the inverse transformation $H^{-1}$. Invertibility of $\mathcal{H}$ enforces a cycle-consistent constraint across frames, which is a regularization for modeling dynamic scenes (Wang et al., 2019).
4. Method
@@
Current approaches for dynamic scene reconstruction typically excel in either 3D geometry reconstruction or novel-view synthesis, but not both. Our framework, illustrated in Fig. 2, introduces a hybrid representation that combines Deformable Gaussian Splatting and Dynamic Neural Surfaces modules, with both components jointly optimized and mutually benefited to enhance performance across both tasks. The details of each module are presented in the following sections.
4.1. DNS for Dynamic Surface Reconstruction
@@
The DNS module utilizes the deformation field defined in Eq. (3) to map the dynamic observation space to canonical space. Unlike other dynamic SDF methods(Cai et al., 2022), our approach incorporates ray-sampling with depth proposals and depth supervision from the DGS module (as shown with orange arrows in fig. 2). Additionally, our method introduces normal supervision using surface normals generated by a foundation model.
@@
Efficient Ray-sampling. Ray-sampling can be computationally intensive without prior information (Mildenhall et al., 2021), as exhaustive sampling is needed to approximate a pixel’s color from an unknown density distribution accurately. Efficiency can be improved by selectively sampling in non-empty regions of the scene. NeRF’s hierarchical volume sampling scheme uses a coarse model to approximate a density distribution, guiding the sampling of the fine model and increasing computational efficiency. In addition to hierarchical sampling, two other major sampling schemes are proposal-based (Barron et al., 2022) and occupancy-based (Yu et al., 2021; Hu et al., 2022). Proposal-based methods, like Mip-NeRF 360 (Barron et al., 2022), replace the coarse model with a compact proposal model that produces only density rather than both density and color. Occupancy-based methods, such as PlenOctree (Yu et al., 2021), effectively filter out points with a low density, avoiding unnecessary sampling. While these coarse-to-fine strategies (Mildenhall et al., 2021; Wang et al., 2021; Rosu and Behnke, 2023; Barron et al., 2022) improve efficiency, they often require costly rendering processes. In our framework, we use the depth map generated by the DGS branch to eliminate unnecessary queries in empty or occluded regions, thereby speeding up the ray-sampling process. The DGS depth map provides proximity to the surface, constraining the sampling range. Specifically, the $\alpha$-blending depth map $d^{\alpha}$ defines these sampling boundaries, and $d^{\alpha}$ is calculated as follows:
@@
where $N$ represents the count of 3D Gaussians encountered, $d_{i}$ is the distance from the $i$-th Gaussian to the camera, and $\alpha_{i}$ denotes opacity. The sampling process starts by emitting a ray from the camera center, $\vec{o}$, along a direction $\vec{v}$. The ray-sampling points in observation space are taken near $\vec{o}+d^{\alpha}\cdot\vec{v}$, with the range adjusted based on SDF values calculated at ray-sampling points, $d=\mathcal{F}(\mathcal{H}(\vec{o}+d^{\alpha}\cdot\vec{v},t))$ , where $\mathcal{F}$ is to predict the SDF value for a point in observation space. The sampling interval along the ray ($\vec{o}$, $\vec{v}$) in observation space is from $\vec{o}+(d^{\alpha}-s|d|)\cdot\vec{v}$ to $\vec{o}+(d^{\alpha}+s|d|)\cdot\vec{v}$, in which the $s$ is a scaling factor for the predicted SDF value $d$.
Figure 3. Visualization of the depth map in 3D space. The leftmost is the RGB image, and images from left to right are 3D point clouds projected from $\alpha$-blending depth, median depth, and filtered depth.
@@
where filtered points are on the surface with SDF loss, as follows:
@@
Apart from depth cues, we introduce the foundation model to provide additional normal supervision to relieve the under-constrained optimization problem. The monocular normal foundation models, i.e. Marigold (Ke et al., 2024; Martin Garcia et al., 2025), are used to generate the pseudo normal map $\bar{{N}}$ . Similar to MonoSDF (Tian et al., 2023), we use the volume rendering method to calculate the normal of ray $\mathbf{r}$, denoted as $\hat{{N}}$ , which is a weighted sum of each ray-sampling point normal, $\nabla_{\mathbf{x_{o}}}\mathcal{F}(\mathcal{H}(x_{o},t)))$ , on the ray $\mathbf{r}$. The consistency between the volume rendered normal map $\hat{{N}}$ and the predicted monocular normal map $\bar{{N}}$ is imposed with angular and $L1$ loss (Tian et al., 2023; Zhang et al., 2024b), as follows: